OpenStudy (anonymous):

The graph shown matches which quadratic equation? A) y = x2 + 3 B) y = x2 - 3 C) y = (x - 3)2 Eliminate D) y = (x + 3)2

OpenStudy (anonymous):

here is the link to the graph . http://usatestprep.com/modules/gallery/files/14/1441/1441.gif

OpenStudy (anonymous):

A

OpenStudy (anonymous):

C) y = (x - 3)2

OpenStudy (king):

A

OpenStudy (anonymous):

A

OpenStudy (anonymous):

for x=0, y=3 therefore, it is A

OpenStudy (anonymous):

or y = x2 + 3

OpenStudy (callisto):

You should know that y=a(x-h)^2 +k where (h,k) is the vertex for your graph, the vertex is at (0,3) so put (h,k) = (0,3) into the equation and neglect the 'a' for your case and you'lll get the answer